OpenStudy (anonymous):

If 832J of energy is required to raise the temperature of a sample of aluminum from 20.0°C to 97.0°C, what mass is the sample of aluminum? (The specific heat of aluminum is 0.90 J/(g × °C).)

OpenStudy (mos1635):

Q=m*c*Δθ Δθ= θfinal-Θ initial

OpenStudy (callisto):

E=mcΔT 832 = m(0.90)(97-20) m=832/[(0.90)(97-20)] = 12.0g